CASE STUDY: Emma is an 18 yr-old freshman at U State where she has received a scholarship to play for the school's soccer team. In high school, Emma dominated as a midfielder, but her U State soccer coach is worried that at her current height of 5'6", weight of 118 lbs, and low body fat of 9% that she needs to increase her body weight before the competitive season or else risk getting injured. To play midfielder within the stop and go, high-intensity sport of soccer, Emma must also possess a high level of endurance. The U State strength coach has set a goal weight for Emma to achieve before the start of the competitive season, of 130 lbs; while maintaining approximately the same low body fat of 9%.

A) What is the MINIMUM amount (grams) of high-quality Protein (PRO) that Emma needs to consume, per dose (i.e., meal, snack, shake, etc), in order to support muscle protein synthesis? 13.4g, 14.8 g, 29.5g, or 32.5 g

B) What is the MINIMUM amount (grams) of high-quality PRO that Emma needs to consume, per meal, in order to support the physical activity demands placed upon her as an athlete and as she works to increase her overall bodyweight? ~15.0 g, ~16.5 g, ~24.7 g, or ~27.1 g

C) On off-season days when there is no practice, what is the MINIMUM amount (grams) of total Carbohydrate (CHO) that Emma needs to consume, per meal, in order to support the physical demands of her sport? ~11.8 g, ~35.4 g, ~59.0 g, or ~70.8 g

D) On off-season days when the team performs 3 hours of practice, what is the MINIMUM amount (grams) of total CHO that Emma needs to consume, per day, in order to support the physical demands of her sport? ~250 g, ~418 g, ~503 g, or ~748 g

E) On off-season days when the team does NOT have practice, what is the MINIMUM amount (kcals) of total Calories that Emma needs to consume, per day, in order to support the physical demands of her sport? ~1856.8 kcal, ~2277.4 kcal, ~2873.4 kcal, or ~4112.5 kcal
Answer #1

answers

here weight is 118 lbs

1 lbs = 0.45 kg

so 118 lbs = 0.45 * 118

=53.1 kg

target weight is 130 lbs = 130* 0.45

= 58.5 kg

1. minimum protien requrement needed per day is 0.8 gm per kilogram body weight

here ,minimum protien needed to intake per day = 53* 0.8

= 42.4 gm protien per day

normaly , we take 3 meals per day , so protien needed from one meal = 42.4 divided by 3

= 14.1 gm per meal

2. in this question approximate amount of protien needed to adjust with extra physical activity is 1.2 gm per kg body weight

= 1.2 *53.1 gm

= 63.72  gm protien per day

= 63.72 / 3

= 21.24 gm per meal

here approximately 24.7 gm protien per meal needed

3. calorie reqirment of a 18 yr old sedentary female is around 1800 per day , in which minimum 40% is contributed by carbohydrats

= 1800 * 40 / 100

= 72,000 / 100

= 720 callorie by carbohydrate

1 gm of carbohydrate provides 4 calorie

so 720 calorie provided by how much gm of carbohydrate = 720 / 4

= 180 gm of carbohydrate per day

= 180 / 3

= 60 gm per meal

4. in moderately excersise 18 yr old female needed around 2000 calorie per day

here minimum 40 % is provided by carhohydratye = 2000* 40 /100

= 80000 / 100

= 800 callorie from CHO

1 gm